Circular No. 6808

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 1997 D. Welch, McMaster University; and W. Liller, Instituto Isaac Newton,

report that a more thorough inspection of the Kodak Technical Pan films taken by Liller in June 1997 (see IAUC 6758) reveals the presence of a definite image at the position of the nova reported on IAUC 6756. The estimated V magnitudes (± 0.3 mag) are as follows: June 5.97 UT, > 13.7; 25.86, 12.7; 30.98, 13.1. The images were revealed by ‘stacking’ the original negatives taken with a 0.20-m f/1.5 Schmidt camera and comparing the position with that of the nova as it appears on the discovery frame taken by the MACHO collaboration.

GRB 980109 K. C. Sahu, Space Telescope Science Institute; and C. Sterken, Uni-

versity of Brussels, report: “The field of GRB 980109 (cf. IAUC 6805) was imaged with the European Southern Observatory’s Dutch 0.91-m telescope at La Silla on Jan. 12.04 and 13.03 UT using a Gunn i filter, and on Jan. 10.11, 11.04, 12.06, and 13.06 using a V filter. The image size is ∼ 3′.7 × 3′.7, and single images centered on the position of the GRB were taken. On Jan. 10 and 11, the observing conditions were poor, the exposure times were 20 min, and the limiting magnitude was ∼ 18.5. These images show no new source when compared with the ESO/Digital Sky Survey fields. On Jan. 12 and 13, the exposure time was 30 min for each image, the sky was photometric, and limiting magnitude of the i images was ∼ 21 (at least 1 mag fainter than the limiting magnitude of the DSS fields). As a result, several new sources are seen, but none of these sources has varied by more than the measurement uncertainty of ∼ 0.2 mag during this 1-day interval. These results indicate that, if the optical counterpart were brighter than I = 20 (as in the case of GRB 970508; cf. IAUC 6654), it lies beyond 1′.5 from the center of the BeppoSAX error box.”

COMET 103P/HARTLEY 2 Visual m_{1} estimates: 1997 Dec. 21.97 UT, 8.2 (J. Bortle, Stormville,

NY, 20×80 binoculars); 26.75, 8.2 (M. Reszelski, Szamotuly, Poland, 20×60 binoculars); 31.07, 8.4 (R. Keen, Mt. Thorodin, CO, 7×50 binoculars); 1998 Jan. 2.55, 8.4 (A. Pearce, Cowaramup, W. Australia, 20×80 binoculars); 11.72, 8.6 (M. Lehký, Hradec Králové, Czech Rep., 25×100 binoculars).

1998 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green